# Quantum Incompatibility in Parallel versus Antiparallel Spins

### v2

@@ Abstract
We explore the joint measurability of incompatible qubit observables on ensembles of parallel and antiparallel spin-½ pairs. In parallel configuration, both spins are prepared in the same state, whereas in antiparallel case, each spin is paired with its flipped counterpart. We show that the antiparallel configuration uniquely enables the exact simultaneous prediction of three mutually orthogonal spin components—an advantage not achievable with parallel states. Extending beyond three observables, we examine joint measurability for larger sets of spin measurements and further generalize our analysis to state configurations beyond the parallel and antiparallel cases. As we show, our results reveal a deep connection to the ‘mean King retrodiction task’ proposed by Vaidman, Aharonov, and Albert, and have implications for a cryptographic protocol introduced by Jeffrey Bub. We further demonstrate how the enhanced compatibility in the antiparallel configuration can facilitate efficient estimation of unknown measurement devices. Finally, we discuss prospects for experimentally realizing the enhanced measurement compatibility in antiparallel configuration by analyzing the effect on finite sub-ensembles of states.
@@
Here, ${\bf a}:=[a_{1},\dots,a_{\mathrm{N}}]\in\{+1,-1\}^{\mathrm{N}}$, where $\mathrm{N}$ is the cardinality of the set $S$, and ${\bf a}\setminus a_{r}$ denotes the summation over all components except $a_{r}$. While sharp spin observables along distinct directions are not jointly measurable, their unsharp versions can be, for suitable values of $\lambda$. For instance, spin observables along the $x$ and $y$ directions are jointly measurable when $\lambda\leq 1/\sqrt{2}$, and observables along $x$, $y$, and $z$ directions are jointly measurable when $\lambda\leq 1/\sqrt{3}$ [30]. For more recent developments, see [32, 33, 34].
Measurement compatibility in multi-copy setting.– More recently, Carmeli et al. [23] have investigated measurement compatibility in a multi-copy setting, where the experimenter has access to multiple copies of a quantum state in each measurement run.
Definition 2.
[Carmeli et al. [23]] A set $S$ of $N$ spin observables is $\mathrm{K}$-copy jointly measurable if there exists a POVM $\mathcal{G}\equiv\{\pi_{{\bf a}}\in\mathcal{P}((\mathbb{C}^{2})^{\otimes\mathrm{K}})\penalty 10000\ |\penalty 10000\ \sum_{{\bf a}}\pi_{{\bf a}}=\mathbf{1}^{\otimes\mathrm{K}}\}$ such that, for all states $\rho_{\vec{m}}$ and all $r,\penalty 10000\ a_{r}$, $\mathrm{Tr}[\rho_{\vec{m}}\,\mathrm{P}^{a_{r}}_{\hat{n}_{r}}(\lambda)]=\sum_{{\bf a}\setminus a_{r}}\mathrm{Tr}[\rho_{\vec{m}}^{\otimes\mathrm{K}}\,\pi_{{\bf a}}]$.
Notably, any set of $N$ spin observables is $\mathrm{K}$-copy jointly measurable whenever $\mathrm{K}\geq N$, by simply performing different measurements on different copies of the state. However, the scenario becomes nontrivial when $\mathrm{K}<N$. As shown in [23], the three mutually unbiased spin observables $X,Y,Z$ (corresponding to directions $\hat{x},\hat{y},\hat{z}$) are 2-copy jointly measurable for sharpness values up to $\lambda=\sqrt{3}/2$, whereas a naive strategy achieves joint measurability only up to $\lambda=1/\sqrt{2}$ (see Fig.1). We now turn to the anti-parallel ($\uparrow\hskip-1.42271pt\downarrow$) configuration, where the measurement is performed on $\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\equiv\rho_{\vec{m}}\otimes\rho_{-\vec{m}}$, with $\rho_{-\vec{m}}:=\tfrac{1}{2}(\mathbf{1}-\vec{m}\cdot\vec{\sigma})$. This leads us to the following definition:
@@
Deferring the proof of Proposition 1 to Supplemental material , here we demonstrate joint measurability of two important sets of spin observables.
@@
Define the set of operators $\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[a,b,c]}\penalty 10000\ |\penalty 10000\ a,b,c=\pm 1\}\subset\mathcal{L}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ as
@@
where $\hskip-1.42271pt[U,V\penalty 10000\ ]\hskip-1.42271pt]:=U\otimes V-V\otimes U$ and $\{\hskip-3.41418pt\{U,V\penalty 10000\ \}\hskip-3.41418pt\}:=U\otimes V+V\otimes U$. Direct computation confirms that each $\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}$ is positive and that $\sum_{a,b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\mathbf{1}^{\otimes 2}$, hence forming a valid POVM $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}\equiv\{\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}\}$. Now, summing over $b$ and $c$ gives $\sum_{b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\tfrac{1}{4}\left(2\mathbf{1}^{\otimes 2}+a\,[\hskip-1.42271pt[X,\mathbf{1}\right]\hskip-1.42271pt])=\mathbb{P}^{a}_{\hat{x}}$. Similarly, summing over $(c,a)$ and $(a,b)$ yields $\mathbb{P}^{b}_{\hat{y}}$ and $\mathbb{P}^{c}_{\hat{z}}$, respectively. Thus, $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}$ satisfies the joint measurability condition of Proposition 1, thereby completing the proof. ∎
It is worth noting that the relative angles between any two observables in the set $\{X,Y,Z\}$ are identical. In the Supplemental Material , we present a detailed analysis showing how this symmetry guides construction of the explicit form of the joint POVM of Eq.(1). There, we also identify a class of other symmetric spin triples that are perfectly jointly measurable in the $2$-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration. In the following, we extend this analysis to examine the joint measurability of four other symmetric spin observables in the $2$-copy $\uparrow\hskip-1.42271pt\downarrow$ configuration.
@@
Consider the set of operators $\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}\penalty 10000\ |\penalty 10000\ a=\pm 1,\ \Gamma\in\{X,Y,Z\}\}\subset\mathcal{L}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$, defined as
@@
$\sum_{\Gamma}\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,+1]}=\tfrac{1}{2}\Bigl(\mathbf{1}^{\otimes 2}+\tfrac{1}{2\sqrt{3}}\sum_{\Gamma}[\hskip-1.42271pt[\Gamma,\mathbf{1}\penalty 10000\ ]\hskip-1.42271pt]\Bigr)=\mathbb{P}^{+1}_{\hat{n}_{0}}.$ (3)
Similarly, summing over the suitable indices reconstructs all $\mathbb{P}^{a_{r}}_{\hat{n}_{r}}$ for $r=0,\cdots 3,\penalty 10000\ \&\penalty 10000\ a_{r}=\pm 1$, satisfying the condition of Proposition 1; and hence completes the proof. ∎
At this point, it is worth recalling a result by Gisin and Popescu [35], which demonstrates that the $\uparrow\hskip-1.42271pt\downarrow$ spin configuration can encode more classical information than the $\uparrow\hskip-1.42271pt\uparrow$ configuration. Remarkably, our Theorems 1 and 2 show that this advantage extends to the joint measurability of multiple spin observables, thereby revealing a novel operational distinction between these configurations.
Recall that a linear map $\Lambda:\mathcal{L}(\mathbb{C}^{2})\to\mathcal{L}(\mathbb{C}^{2})$ is said to be positive (P) if it maps density operators to density operators, i.e., $\Lambda:\mathcal{D}(\mathbb{C}^{2})\to\mathcal{D}(\mathbb{C}^{2})$. However, physicality further requires complete positivity (CP), meaning that the extended map $\mathrm{id}_{d}\otimes\Lambda$ must also preserve positivity on $\mathcal{D}(\mathbb{C}^{d}\otimes\mathbb{C}^{2})$ for all $d\geq 2$, where $\mathrm{id}_{d}$ is the identity map on $\mathbb{C}^{d}$. Thanks to the Choi–Jamiołkowski isomorphism [26, 27], it suffices to verify complete positivity using $d=2$ for qubit maps. In this work, we focus on trace-preserving (TP) maps, which satisfy $\operatorname{Tr}[\Lambda(A)]=\operatorname{Tr}[A]$ for all $A\in\mathcal{L}(\mathbb{C}^{2})$. The spin-flip map $\mathrm{F}$, which appears in the $\uparrow\hskip-1.42271pt\downarrow$ configuration, is an example of a PTP map that is not CPTP (PnCP). Moreover, it is self dual, namely $\mathrm{F}=\mathrm{F}^{\star}$ [36], and $\mathrm{F}\circ\mathrm{F}=\mathrm{id}_{2}$, where $\circ$ denotes sequential composition. This observation motivates us to go beyond the specific $\uparrow\hskip-1.42271pt\uparrow$ and $\uparrow\hskip-1.42271pt\downarrow$ configurations and explore joint measurability in more general configurations of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is either a CPTP or a PTP map. At first glance, the question of joint measurability on a configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ may seem ill-posed when $\Lambda$ is PnCP, as such maps are not physically implementable. However, the question becomes meaningful from an adversarial perspective: a Referee, who knows the state, prepares the bipartite configuration and hands it to an experimentalist, who is unaware of the preparation but aims to implement a joint measurement. Later, we will illustrate how such configurations can arise naturally in various practical scenarios. Before that, we first establish a general result on joint measurability when such broader settings are considered (proof provided in End Matter).
@@
Theorem 3 shows that to achieve an advantage in joint measurability over the $\uparrow\hskip-1.42271pt\uparrow$ configuration, one must consider configurations of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is PnCP. A pertinent example is provided by the family of maps $\mathrm{F}_{\mu}(\rho_{\vec{m}}):=\tfrac{1}{2}(\mathbf{1}-\mu\,\vec{m}\cdot\vec{\sigma})$, which interpolate between the identity map and the spin-flip map, i.e., $\mathrm{F}_{\mu}=\tfrac{1-\mu}{2}\,\text{id}_{2}+\tfrac{1+\mu}{2}\,\mathrm{F},\quad\mu\in[-1,1]$. These maps are PnCP for $\mu\in(1/3,1]$, but CPTP otherwise [37, 38, 39].
@@
The configuration $\rho_{\vec{m}}\otimes\mathrm{F}_{\mu}(\rho_{\vec{m}})$ offers an enhanced joint measurability of unsharp MUB observables compared to the $\uparrow\hskip-1.42271pt\uparrow$ configuration whenever $\mu>(\sqrt{3}-1)$.
Proof is discussed in the End Matter. Corollary 1, however, does not demonstrate an advantage over the $\uparrow\hskip-1.42271pt\uparrow$ configuration in the intermediate regime $\mu\in(1/3,\sqrt{3}-1]$, where $\mathrm{F}_{\mu}$ is PnCP but the joint measurability enhancement remains unproven for MUB observables. Whether an advantage exists in this regime for other sets of observables remains open for further investigation. As we discuss in Supplemental , the enhanced performance observed in the $\uparrow\hskip-1.42271pt\downarrow$ case (and in the configuration considered in Corollary 1) can be better appreciated from the mathematical structure of generalized probabilistic theories (GPTs) [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. In the End Matter we introduce the notion of joint measurability for a set of spin observables restricted to a sub-ensemble of qubit states $\mathcal{D}_{SE}\subset\mathcal{D}(\mathbb{C}^{2})$, which provide an experimental means to validate the advantages established in Theorems 1, 2, and Corollary 1. In the remainder of the manuscript, we explore connections to other foundational concepts and discuss the potential applications.
@@
Building on VAA’s mean King strategy, Bub proposed a novel quantum key distribution (QKD) protocol in which Bob restricts his measurements to two choices (e.g., $X$ and $Z$), and Alice performs the $\mathbb{M}_{+}$ measurement on her ancilla and the channel particle returned by Bob [29] (see also [57]). To detect eavesdropping, Alice publicly announces her retrodicted outcomes whenever the projectors $\xi_{[+1,-1,-1]}$ or $\xi_{[-1,-1,+1]}$ click. Bob then checks whether her retrodiction matches his actual result. If no discrepancies arise—indicating no interference by Eve—they use the remaining outcomes to generate a shared raw key (see Table 1). Unlike the BB84 [58] and E91 [59] protocols, this scheme allows key generation without revealing measurement bases, provided Bob’s choices are limited to two observables. However, when Bob chooses among all three Pauli observables ($X$, $Y$, and $Z$), basis disclosure becomes necessary. Interestingly, as shown in Table 1, the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ supports secure key generation even in this case.
Consider the scenario where Bob performs unsharp spin measurements. As shown in [30], the unsharp observables $\{X(\lambda),Z(\lambda)\}$ are jointly measurable for $\lambda\leq 1/\sqrt{2}$, while the triple $\{X(\lambda),Y(\lambda),Z(\lambda)\}$ is compatible for $\lambda\leq 1/\sqrt{3}$. Being compatibility within $\lambda\in[0,1/\sqrt{2}]$, the original protocol yields zero key rate, whereas the modified scheme employing three unsharp observables has the potential to offer an advantage for $\lambda\in(1/\sqrt{3},1/\sqrt{2}]$. While a detailed analysis of this intuition is left for future work, we now turn to another potential application of measurement compatibility in the $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Estimating unknown measurement devices.– Estimation of unknown quantum resources—such as states, channels, or measurement devices—is essential for the reliable implementation of quantum protocols [60, 61, 62, 63, 64, 65]. When the resource is supplied by an untrusted party, a Verifier seeks to verify its identity as efficiently as possible. Here, we consider the task of estimating an unknown qubit projective measurement $\mathrm{M}_{\hat{n}}\equiv\{\mathrm{P}^{a}_{\hat{n}}\penalty 10000\ |\penalty 10000\ a=\pm 1\}$, where the measurement axis $\hat{n}$ is uniformly distributed over the Bloch sphere. The Verifier treats the device as a black box with trusted outcome labels and aims to estimate $\hat{n}$ with minimal use of the device.
@@
Our study welcomes several questions for future investigation. While Theorems 1 and 2 demonstrate the advantage of the antiparallel configuration for specific choices of spin observables, it remains an open question whether similar advantages persist for more general sets of observables. Furthermore, Corollary 1 establishes an advantage of the configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ over the parallel configuration for a certain class of PnCP maps. It would be interesting to explore whether this advantage extends to all PnCP maps when considering more general observables. Finally, extending the analysis to higher-dimensional quantum systems may reveal additional structures and potential applications.

### v3

@@ Abstract
We explore the joint measurability of incompatible qubit observables on ensembles of parallel and antiparallel spin-½ pairs. In parallel configuration, both spins are prepared in the same state, whereas in antiparallel case, each spin is paired with its flipped counterpart. We show that the antiparallel configuration uniquely enables exact simultaneous prediction of three mutually orthogonal spin components—an advantage not achievable with parallel states. Extending beyond three observables, we examine joint measurability for larger sets of spin measurements and further generalize our analysis to state configurations beyond the parallel and antiparallel cases. As we show, our results reveal a deep connection to the ‘mean King retrodiction task’ proposed by Vaidman, Aharonov, and Albert, and have implications for a cryptographic protocol introduced by Jeffrey Bub. We further demonstrate how the enhanced compatibility in the antiparallel configuration can facilitate efficient estimation of unknown measurement devices. Finally, we discuss prospects for experimentally realizing the enhanced measurement compatibility in antiparallel configuration by analyzing the effect on finite sub-ensembles of states.
@@
Here, ${\bf a}:=[a_{1},\dots,a_{\mathrm{N}}]\in\{+1,-1\}^{\mathrm{N}}$, where $\mathrm{N}$ is the cardinality of the set $S$, and ${\bf a}\setminus a_{r}$ denotes the summation over all components except $a_{r}$. While sharp spin observables along distinct directions are not jointly measurable, their unsharp versions can be, for suitable values of $\lambda$. For instance, spin observables along the $x$ and $y$ directions are jointly measurable when $\lambda\leq 1/\sqrt{2}$, and observables along $x$, $y$, and $z$ directions are jointly measurable when $\lambda\leq 1/\sqrt{3}$ [30] (see also [32, 33, 34]).
Measurement compatibility in multi-copy setting.– More recently, Carmeli et al. [23] have investigated measurement compatibility in a multi-copy setting, where the experimenter has access to multiple copies of a quantum state in each measurement run. Notably, any set of $N$ spin observables is $\mathrm{K}$-copy jointly measurable whenever $\mathrm{K}\geq N$, by simply performing different measurements on different copies of the state. However, the scenario becomes nontrivial when $\mathrm{K}<N$. As shown in [23], the three mutually unbiased spin observables $X,Y,Z$ (corresponding to directions $\hat{x},\hat{y},\hat{z}$) are 2-copy jointly measurable for sharpness values up to $\lambda=\sqrt{3}/2$, whereas a naive strategy achieves joint measurability only up to $\lambda=1/\sqrt{2}$ (see Fig.1). We now turn to the anti-parallel ($\uparrow\hskip-1.42271pt\downarrow$) configuration, where the measurement is performed on $\rho_{\vec{m}}\otimes\mathrm{F}(\rho_{\vec{m}})\equiv\rho_{\vec{m}}\otimes\rho_{-\vec{m}}$, with $\rho_{-\vec{m}}:=\tfrac{1}{2}(\mathbf{1}-\vec{m}\cdot\vec{\sigma})$. This leads us to the following definition:
@@
Deferring the proof of Proposition 1 to Supplemental material [35], here we demonstrate joint measurability of two important sets of spin observables.
@@
(Outline) A straightforward calculation confirms the set of operators
@@
to be positive for $\penalty 10000\ a,b,c=\pm 1$; here $\hskip-1.42271pt[U,V\penalty 10000\ ]\hskip-1.42271pt]:=U\otimes V-V\otimes U$ and $\{\hskip-3.41418pt\{U,V\penalty 10000\ \}\hskip-3.41418pt\}:=U\otimes V+V\otimes U$. Furthermore, $\sum_{a,b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\mathbf{1}^{\otimes 2}$, ensure $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}\equiv\big\{\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}\penalty 10000\ |\penalty 10000\ a,b,c=\pm 1\big\}$ to be a POVM on $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. Summing over $b$ and $c$ yields $\sum_{b,c}\Pi_{[a,b,c]}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}=\tfrac{1}{4}\left(2\mathbf{1}^{\otimes 2}+a\,[\hskip-1.42271pt[X,\mathbf{1}\right]\hskip-1.42271pt])=\mathbb{P}^{a}_{\hat{x}}$. Similarly, summing over $(c,a)$ and $(a,b)$ yields $\mathbb{P}^{b}_{\hat{y}}$ and $\mathbb{P}^{c}_{\hat{z}}$, respectively. Thus, $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{MUB}}$ satisfies the joint measurability condition of Proposition 1, thereby completing the proof. ∎
As discussed in the End Matter, symmetries play a crucial role in the explicit construction of the joint POVM in Theorem 1. In the Supplemental Material [35], we generalize this construction to a class of three symmetric spin observables. In the next, we rather present joint measurability of four symmetric spin observables on antiparallel configuration.
@@
(Outline) The set of operators $\mathcal{G}^{\uparrow\hskip-1.42271pt\downarrow}_{\mathrm{SIC}_{4}}\equiv\Big\{\Pi^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{[\Gamma,a]}\penalty 10000\ |\penalty 10000\ a=\pm 1,\ \Gamma\in\{X,Y,Z\}\Big\}$, where
@@
We recall here the result of Gisin and Popescu [36] which reports that the $\uparrow\hskip-1.42271pt\downarrow$ spin configuration encodes more classical information than the $\uparrow\hskip-1.42271pt\uparrow$ configuration. Theorems 1 and 2 show that this advantage extends to the joint measurability of multiple spin observables. Using semidefinite programming, we further establish that this advantage is generic for any triple of spin observables, except when they lie in a common plane (see End Matter).
The spin-flip map $\mathrm{F}$ appearing in the $\uparrow\hskip-1.42271pt\downarrow$ configuration is a positive trace-preserving (PTP) but not completely positive (CPTP) map [37], which we refer to as PnCP. Moreover, it is self-dual, $\mathrm{F}=\mathrm{F}^{\star}$ [38], and involutive, $\mathrm{F}\circ\mathrm{F}=\mathrm{id}_{2}$. These properties motivate us to go beyond the specific $\uparrow\hskip-1.42271pt\uparrow$ and $\uparrow\hskip-1.42271pt\downarrow$ configurations and study joint measurability for bipartite states of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$, where $\Lambda$ is either CPTP or PTP. Although joint measurability with PnCP maps may appear ill-posed, it is operationally meaningful in an adversarial setting: a Referee, knowing the state, prepares the configuration and hands it to an experimentalist who is unaware of the preparation but attempts a joint measurement. As we will discuss later, such configurations arise naturally in several practical settings. We first establish a general result on joint measurability in this broader framework; its proof is deferred to the End Matter.
@@
An advantage in joint measurability over the $\uparrow\hskip-1.42271pt\uparrow$ configuration, thus, arises only for states of the form $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ with $\Lambda$ PnCP. A representative family is $\mathrm{F}_{\mu}(\rho_{\vec{m}})=\tfrac{1}{2}(\mathbf{1}-\mu\vec{m}\cdot\vec{\sigma})=\tfrac{1-\mu}{2}\mathbf{1}_{2}+\tfrac{1+\mu}{2}\mathrm{F},\quad\mu\in[-1,1]$, which is PnCP for $\mu\in(1/3,1]$ and CPTP otherwise [39, 40, 41].
@@
The configuration $\rho_{\vec{m}}\otimes\mathrm{F}_{\mu}(\rho_{\vec{m}})$ outperforms the $\uparrow\hskip-1.42271pt\uparrow$ configuration in the joint measurability of unsharp MUB observables whenever $\mu>\sqrt{3}-1$.
The proof is deferred to the End Matter. We emphasize, however, that PnCP maps do not generically guarantee an advantage as in Corollary 1. For example, consider the PnCP pancake map $\Lambda^{\mathrm{PC}}(\rho_{\vec{m}}):=\tfrac{1}{2}(\rho_{\vec{m}}+\rho_{\vec{m}}^{\mathrm{T}})=\tfrac{1}{2}(\mathbf{1}+m_{x}X+m_{z}Z)$, where $\mathrm{T}$ denotes transposition in the computational basis.
Observation 1.
The spin observables $Y$ and $(Z+Y)/\sqrt{2}$ are jointly measurable on the $\uparrow\hskip-1.42271pt\uparrow$ configuration, but not on $\rho_{\vec{m}}\otimes\Lambda^{\mathrm{PC}}(\rho_{\vec{m}})$.
Joint measurability on the $\uparrow\hskip-1.42271pt\uparrow$ configuration is trivial. In contrast, for the eigenstates of $Y\equiv\rho_{Y}^{\pm}=\tfrac{1}{2}(\mathbf{1}\pm Y)$ one has $\rho_{Y}^{\pm}\otimes\Lambda^{\mathrm{PC}}(\rho_{Y}^{\pm})=\tfrac{1}{2}(\mathbf{1}\pm Y)\otimes\tfrac{\mathbf{1}}{2}$, which shows that a sharp $Y$ statistics can be obtained only if $Y$ is performed on the first copy, while the second copy lacks sharp statistics for $(Z+Y)/\sqrt{2}$. Thus whether $\mathrm{F}_{\mu}$ yields an advantage in the intermediate regime $\mu\in(1/3,\sqrt{3}-1]$ for other choices of observables remains an open question.
The enhanced measurement compatibility in the $\uparrow\hskip-1.42271pt\downarrow$ configuration (and in Corollary 1) can be better appreciated within the framework of generalized probabilistic theories (GPTs) [42, 43, 44, 45, 46, 47, 48, 49]. Assuming local systems are quantum, GPTs allow multiple prescriptions for composing subsystems, interpolating between minimal and maximal tensor products [50, 51, 52, 53, 54, 55]. The minimal tensor product—containing only separable states but an enlarged set of effects—permits composite measurements forbidden in standard quantum theory, rendering certain incompatible observables jointly measurable. In particular, we establish the following result.
Theorem 4.
For any set of qubit POVMs, joint measurability on the $\uparrow\hskip-1.42271pt\downarrow$ configuration in quantum theory implies joint measurability on the $\uparrow\hskip-1.42271pt\uparrow$ configuration in minimal tensor product composition.
The proof is deferred to the Supplemental Material, and there we also note that the converse statement of Theorem 4 does not hold in general. In the End Matter, we introduce joint measurability restricted to a sub-ensemble $\mathcal{D}_{\mathrm{SE}}\subset\mathcal{D}(\mathbb{C}^{2})$, providing an experimentally accessible route to test the advantages established in Theorems 1, 2, and Corollary 1. The remainder of the manuscript shows connections to foundational concepts and potential applications.
@@
Building on the mean King strategy of VAA, Bub proposed a quantum key distribution (QKD) protocol in which Bob restricts his measurements to two observables (e.g., $X$ and $Z$), while Alice performs the POVM $\mathbb{M}_{+}$ jointly on her ancilla and the particle returned by Bob [29] (see also [60]). To detect eavesdropping, Alice publicly announces her retrodicted outcomes whenever the projectors $\xi_{[+1,-1,-1]}$ or $\xi_{[-1,-1,+1]}$ occur. Bob then verifies whether her retrodiction agrees with his actual measurement outcome. In the absence of discrepancies—indicating no intervention by an eavesdropper—the remaining outcomes are used to generate a shared raw key (see Table 1). Unlike the BB84 [61] and E91 [62] protocols, this scheme enables key generation without revealing measurement bases, provided Bob’s choices are restricted to two observables. If Bob instead chooses among all three Pauli observables $X,Y,$ and $Z$, basis disclosure is generally required. Remarkably, as shown in Table 1, the POVM $\mathcal{G}^{{}_{\uparrow\hskip-1.42271pt\downarrow}}_{\mathrm{MUB}}$ enables secure key generation even in this three-observable setting.
We now consider the case where Bob performs unsharp spin measurements. As shown in [30], the pair $\{X(\lambda),Z(\lambda)\}$ is jointly measurable for $\lambda\leq 1/\sqrt{2}$, whereas the triple $\{X(\lambda),Y(\lambda),Z(\lambda)\}$ is jointly measurable only for $\lambda\leq 1/\sqrt{3}$. Consequently, for $\lambda\in[0,1/\sqrt{2}]$, the original two-observable protocol lies entirely within a compatibility regime and therefore yields no key. In contrast, the modified scheme employing three unsharp observables accesses the intermediate regime $\lambda\in(1/\sqrt{3},1/\sqrt{2}]$, where pairwise incompatibility persists while triple-wise compatibility is absent. This identifies a parameter range in which the enhanced compatibility structure enabled by the $\uparrow\hskip-1.42271pt\downarrow$ configuration promises improved noise robustness to implement Bub’s protocol. A detailed analysis of key-rate calculations and noise-tolerance bounds is beyond the scope of the present Letter and is left for future work. We rather turn to another potential application of enhanced measurement compatibility in the $\uparrow\hskip-1.42271pt\downarrow$ configuration.
Estimating unknown measurement devices.– Estimation of unknown quantum resources—such as states, channels, or measurement devices—is a central task for the reliable implementation of quantum protocols [63, 64, 65, 66, 67, 68]. When a resource is supplied by an untrusted party, a verifier seeks to identify it as efficiently as possible using minimal experimental resources. Here, we consider the task of estimating an unknown qubit projective measurement $\mathrm{M}_{\hat{n}}\equiv\{\mathrm{P}^{a}_{\hat{n}}\mid a=\pm 1\}$, where the measurement axis $\hat{n}$ is uniformly distributed over the Bloch sphere. The verifier treats the device as a black box with trusted outcome labels and aims to estimate $\hat{n}$ using as few invocations of the device as possible.
[…]
@@
Our study welcomes several questions for future investigation. While Theorems 1 and 2 demonstrate the advantage of the antiparallel configuration for specific choices of spin observables, it remains an open question whether similar advantages persist for more general sets of observables. Furthermore, Corollary 1 establishes an advantage of the configuration $\rho_{\vec{m}}\otimes\Lambda(\rho_{\vec{m}})$ over the parallel configuration for a certain class of PnCP maps. It would be interesting to explore for which class of PnCP maps such an advantage will be possible. Finally, extending the analysis to higher-dimensional quantum systems may reveal additional structures and potential applications.